# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v1

@@ 3.2 Theoretical Analysis
for which
$\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})={\bf f}(\widehat{\bm{\theta}},{\bf x})+{{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}\left({\bf w}-{\mathbf{J}}_{\mathcal{X}}\widehat{\bm{\theta}}\right)$
$+\left({{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}-{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x})\right){\bf z}_{0},$ (6)
where ${{\mathbf{K}}}_{{\bf x},\mathcal{X}}\triangleq{{\mathbf{K}}}_{\widehat{\bm{\theta}}}({\bf x},\mathcal{X})\in^{c\times nc}$. Taking ${\bf z}_{0}$ to be a random variable, we form an ensemble of predictors $\{\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\}_{{\bf z}}$, where each $\bm{\theta}^{\star}$ is formed from the projection of a random ${\bf z}$ onto $\text{Null}({\mathbf{J}}_{\mathcal{X}})$. We require ${\bf z}$’s distribution to be symmetric, isotropic, and centered at $\widehat{\bm{\theta}}$, as we do not know a priori which directions contain more information. We choose the maximum entropy distribution for a given mean and variance, which is Gaussian [^1] Heavier-tailed distributions matching the above criteria, e.g. logistic distributions, may improve results.. Hence, we let ${\bf z}_{0}\sim\mathcal{N}({\bf 0},\gamma^{2}\mathbf{I})$, for some hyper-parameter $\gamma\in$. The expectation and variance of Section 3.2 are ${\mathbb{E}}\big{(}\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\big{)}=\bm{\mu}(\widehat{\bm{\theta}},{\bf x}),\text{Var}\big{(}\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\big{)}=\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})$, where
$\mathbf{\bm{\mu}}(\widehat{\bm{\theta}},{\bf x})$ $={{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}(\bm{\theta}^{\ddagger}-\widehat{\bm{\theta}})+{\bf f}(\widehat{\bm{\theta}},{\bf x}),$ (7)
$\mathbf{\bm{\sigma}}^{2}(\widehat{\bm{\theta}},{\bf x})$ $=\big{(}{{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big{)}\gamma^{2}.$ (8)
The distribution of the predictor ${\bf f}(\bm{\theta}^{\star},{\bf x})$ is then
[…]
@@
$\begin{bmatrix}{\bf f}(\bm{\theta}^{\star},\mathcal{X})\\ f(\bm{\theta}^{\star},{\bf x})\end{bmatrix}\!$ $\sim\mathcal{N}\Bigg{(}\!\!\begin{bmatrix}{\bf f}(\widehat{\bm{\theta}},\mathcal{X})\\ f(\widehat{\bm{\theta}},{\bf x})\end{bmatrix}\!,\gamma^{2}\!\begin{bmatrix}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}&{{\bf k}}_{{\bf x},\mathcal{X}}\\ {{\bf k}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}&\kappa_{{\bf x},{\bf x}}\end{bmatrix}\!\!\Bigg{)}.\!\!$ (10)
@@
Beyond quadratic loss, for a general loss function satisfying the assumptions of Lemma 3.1, a clear GP posterior interpretation like that in 10 may not exist. Nevertheless, we can still derive related insights. If $\widehat{\bm{\theta}}$ is an interpolating solution from initial training, which is common for modern NNs, then as long as ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank, solving 3 is equivalent to finding $\bm{\theta}\in^{p}$ such that $\widetilde{{\bf f}}(\bm{\theta},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., find $\bm{\theta}\in^{p}$ such that ${\mathbf{J}}_{\mathcal{X}}(\bm{\theta}-\widehat{\bm{\theta}})={\bf 0}$. So we obtain (9) with $\bm{\mu}(\widehat{\bm{\theta}},{\bf x})={\bf f}(\widehat{\bm{\theta}},{\bf x})$ and $\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})=\big{(}{{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big{)}\gamma^{2}$ as the conditional distribution of 10, conditioned on the event ${\bf f}(\bm{\theta}^{\star},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., interpolation with $\ell({\bf f}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})=0$ for $i=1,\ldots,n$.
@@
Drawing samples from the posterior 9 by explicitly calculating 7 and 8 can be intractable in large-scale settings. Moreover, it can be numerically unstable due to the highly ill-conditioned nature of the NTK matrix [^2] Recall that the condition number of ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}$ is the square of that of ${\mathbf{J}}_{\mathcal{X}}$.. However, by combining the above derivations with Theorem 3.2, we arrive at the key property of Algorithm 1: it enables efficient sampling from the posterior 9.
@@
With the assumptions of Theorem 3.2, the samples generated by Algorithm 1 represent draws from the predictive distribution in 9.
Hence, we approximate 7 and 8 by computing the sample mean and covariance of $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x})\}_{s=1}^{S}$ obtained from Algorithm 1. By the law of large numbers, the quality of these approximations improves as $S\to\infty$.
@@
Loss functions that do not satisfy the assumption of Lemma 3.1, such as the cross-entropy loss, may fail to yield a unique representation of 4, so the above posterior analysis does not apply. However, our experiments demonstrate that Algorithm 1 can still generate samples that effectively capture the posterior variance (see Section 5.3) and posterior mean (see Section 5.4). In cases where 3 lacks a solution, Algorithm 1 can still be executed by terminating the iterations of (stochastic) GD early. Investigating the distribution of the resulting ensemble and its connection to an explicit posterior remains a potential direction for future research.
4 Related Works
NUQLS shares notable similarities with, and exhibits key differences from, several prior works that we briefly allude to here; please see Appendix B for more discussions.
LLA Framework. Like the popular LLA framework (Khan et al., 2019; Foong et al., 2019; Immer et al., 2021; Daxberger et al., 2021), NUQLS relies on linearizing the network, but it diverges in its approach to posterior approximation. While LLA requires sampling from a parameter distribution and often imposes a prior to handle degeneracy in overparameterized settings, NUQLS directly approximates the predictive distribution without artificial priors, avoiding biases and additional hyperparameters. This results in a noise-free GP with an NTK kernel, making NUQLS more suitable for interpolating regimes (Hodgkinson et al., 2023). Additionally, NUQLS captures output covariances in classification tasks, offering a more comprehensive predictive distribution than LLA, which produces independent GPs for each output. NUQLS also provides flexibility in regression tasks by allowing post-hoc variance scaling (see $\gamma$ in 9), enabling efficient hyperparameter tuning without retraining (see Section C.1 for an explanation of this method).
Among the LLA-inspired methods, Sampling-LLA in Antorán et al. (2022) bears the closest resemblance to NUQLS. However, in addition to the fundamental differences between NUQLS and LLA-inspired methods mentioned above, Sampling-LLA has several distinctions from NUQLS. Above all, unlike Sampling-LLA, NUQLS avoids the need for regularization and random subproblems. Instead, NUQLS leverages the true training data and random initialization to naturally capture uncertainty in overparameterized settings. Another related approach is Variational LLA (Ortega et al., 2023), which uses a variational sparse GP with an NTK to approximate the LLA predictive distribution. Variational LLA has been shown to outperform Sampling-LLA in terms of both computational efficiency and uncertainty quantification performance.
Ensemble Framework. NUQLS also differs from ensemble-based methods. Unlike Bayesian Deep Ensembles (BDE) in (He et al., 2020), which modifies infinitely wide networks to derive a GP predictive distribution, NUQLS demonstrates that finite-width, unmodified linearized networks inherently sample from a GP with an NTK kernel. Similarly, the local ensembles method (Madras et al., 2019), which perturbs the NN along small-curvature directions to create nearly loss-invariant ensembles, differs from NUQLS in that it relies on second-order information and includes directions with small but non-zero curvature. NUQLS, by contrast, constructs ensembles with exactly the same loss level from using only first-order information, providing a more precise representation of uncertainty. Low-rank approximations used in local ensembles may also poorly capture the true curvature structure, further distinguishing the two approaches.
@@ 4 Related Works
Table 1: Comparing performance of NUQLS, DE, LLA and SWAG on UCI regression tasks. NUQLS performs as well as or better than all other methods, while showing a speed up over other methods; this speed up increases with the size of the datasets. LLA-K denotes LLA with a KFAC covariance structure.
@@
Energy NUQLS $\mathbf{0.047\pm}$$\mathbf{0.006}$ $\mathbf{-2.400\pm}$$\mathbf{0.209}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $\mathbf{0.151}$
DE $0.218\pm$$0.032$ $\mathbf{-1.651\pm}$$\mathbf{0.783}$ $\mathbf{0.004\pm}$$\mathbf{0.002}$ $102.244$
LLA $\mathbf{0.048\pm}$$\mathbf{0.006}$ $\mathbf{-2.475\pm}$$\mathbf{0.128}$ $\mathbf{0.004\pm}$$\mathbf{0.004}$ $0.269$
SWAG $\mathbf{0.058\pm}$$\mathbf{0.015}$ $-1.950\pm$$0.158$ $0.080\pm$$0.011$ $37.084$
Kin8nm NUQLS $\mathbf{0.252\pm}$$\mathbf{0.005}$ $-0.796\pm$$0.025$ $\mathbf{0.000\pm}$$\mathbf{0.000}$ $\mathbf{0.264}$
DE $\mathbf{0.252\pm}$$\mathbf{0.006}$ $\mathbf{-0.914\pm}$$\mathbf{0.028}$ $0.002\pm$$0.001$ $73.967$
LLA $\mathbf{0.260\pm}$$\mathbf{0.010}$ $-0.783\pm$$0.054$ $\mathbf{0.001\pm}$$\mathbf{0.001}$ $11.966$
SWAG $0.457\pm$$0.149$ $-0.006\pm$$0.295$ $0.054\pm$$0.012$ $150.263$
Protein NUQLS $\mathbf{0.623\pm}$$\mathbf{0.005}$ $\mathbf{0.209\pm}$$\mathbf{0.047}$ $\mathbf{0.002\pm}$$\mathbf{0.000}$ $\mathbf{1.356}$
DE $0.741\pm$$0.052$ $\mathbf{0.203\pm}$$\mathbf{0.203}$ $\mathbf{0.011\pm}$$\mathbf{0.020}$ $1014.827$
LLA-K $0.640\pm$$0.007$ $0.458\pm$$0.071$ $\mathbf{0.002\pm}$$\mathbf{0.000}$ $9.506$
SWAG $0.730\pm$$0.044$ $\mathbf{0.187\pm}$$\mathbf{0.080}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $468.972$
Figure 2: Violin plot of VMSP, for correctly predicted ID test points, incorrectly predicted ID test points, and OoD test points. Median is shown, with violin width depicting density. Low variance is expected for ID correct points, and large variance for ID incorrect and OoD points. Here FMNIST denotes FashionMNIST, MC denotes MC-Dropout, and BASE denotes Baseline.
Table 2: Image classification predictive performance, using LeNet5 on MNIST and FashionMNIST (FMNIST). Experiment was run $5$ times with different random MAP initialisations to get standard deviation on metrics.
MAP $0.034\pm$$0.002$ $\mathbf{0.990\pm}$$\mathbf{0.001}$ $0.008\pm$$0.001$ $0.888\pm$$0.008$ $0.886\pm$$0.008$ $257$
NUQLS $0.035\pm$$0.002$ $0.989\pm$$0.001$ $\mathbf{0.003\pm}$$\mathbf{0.000}$ $\mathbf{0.930\pm}$$\mathbf{0.026}$ $\mathbf{0.928\pm}$$\mathbf{0.026}$ $106$
MNIST DE $0.034\pm$$0.004$ $\mathbf{0.991\pm}$$\mathbf{0.000}$ $0.011\pm$$0.004$ $\mathbf{0.932\pm}$$\mathbf{0.009}$ $\mathbf{0.928\pm}$$\mathbf{0.009}$ $2845$
MC-Dropout $0.044\pm$$0.002$ $0.989\pm$$0.000$ $0.017\pm$$0.01$ $0.873\pm$$0.032$ $0.871\pm$$0.031$ $533$
SWAG $\mathbf{0.029\pm}$$\mathbf{0.003}$ $\mathbf{0.991\pm}$$\mathbf{0.000}$ $\mathbf{0.004\pm}$$\mathbf{0.002}$ $0.902\pm$$0.008$ $0.900\pm$$0.008$ $489$
LLA* $0.034\pm$$0.002$ $\mathbf{0.990\pm}$$\mathbf{0.001}$ $0.008\pm$$0.001$ $0.888\pm$$0.008$ $0.886\pm$$0.008$ $45$
VaLLA $0.034\pm$$0.002$ $\mathbf{0.990\pm}$$\mathbf{0.001}$ $0.008\pm$$0.001$ $0.889\pm$$0.008$ $0.886\pm$$0.008$ $1583$
MAP $0.298\pm$$0.007$ $0.891\pm$$0.3$ $\mathbf{0.006\pm}$$\mathbf{0.001}$ $0.840\pm$$0.022$ $0.804\pm$$0.021$ $158$
NUQLS $0.302\pm$$0.006$ $0.891\pm$$0.002$ $\mathbf{0.005\pm}$$\mathbf{0.002}$ $\mathbf{0.904\pm}$$\mathbf{0.007}$ $\mathbf{0.870\pm}$$\mathbf{0.006}$ $89$
FMNIST DE $0.288\pm$$0.002$ $0.896\pm$$0.001$ $0.013\pm$$0.001$ $0.876\pm$$0.003$ $0.836\pm$$0.003$ $1587$
MC-Dropout $0.306\pm$$0.007$ $0.892\pm$$0.003$ $0.026\pm$$0.002$ $0.856\pm$$0.021$ $0.813\pm$$0.019$ $291$
SWAG $\mathbf{0.283\pm}$$\mathbf{0.005}$ $\mathbf{0.899\pm}$$\mathbf{0.003}$ $0.018\pm$$0.002$ $0.817\pm$$0.023$ $0.783\pm$$0.022$ $264$
LLA* $0.298\pm$$0.007$ $0.891\pm$$0.003$ $\mathbf{0.006\pm}$$\mathbf{0.001}$ $0.841\pm$$0.022$ $0.805\pm$$0.021$ $26$
VaLLA $0.298\pm$$0.007$ $0.891\pm$$0.003$ $0.007\pm$$0.001$ $0.841\pm$$0.022$ $0.805\pm$$0.021$ $1583$
5 Experiments
We now compare the performance of our method with alternatives on various regression and classification tasks. Implementation details are given in Appendix E. The PyTorch implementation of our experiments is available here.
5.1 Toy Regression
We compare the performance of our method on a toy regression problem, taken from (Hernández-Lobato & Adams, 2015) and extended in (Park & Blei, 2024). In Figure 1, we take $20$ uniformly sampled points in the domain $x\in[-4,-2]\cup[2,4]$, and let $y=x^{3}+\epsilon,~{}\epsilon\sim\mathcal{N}(0,3^{2})$. A small MLP was trained on these data and used for prediction. We apply VI, SWAG, LA, LLA, DE and NUQLS to the network to find a predictive mean and uncertainty. Close to the training data, i.e. in $[-4,-2]\cup[2,4]$ we expect low uncertainty; outside of this region, the uncertainty should grow with distance from the training points. VI underestimates, while SWAG and LA overestimate the uncertainty. DE grows more uncertain with distance from the training points, however both NUQLS and the LLA contain the underlying target curve within their confidence intervals. Note that deep ensembles output a heteroskedastic variance term, and were trained on a Gaussian likelihood; in comparison, the variances for LLA and NUQLS were computed post-hoc.
5.2 UCI Regression
@@
In Tables 1 and 6, we compare NUQLS with DE, LLA and SWAG on a series of UCI regression problems. Mean squared error (MSE) and expected calibration error (ECE) respectively evaluate the predictive and UQ performance, with Gaussian negative log likelihood (NLL) evaluating both. See (Nemani et al., 2023a, §$\S 4.11$) for an explanation of ECE. We see that NUQLS consistently has the (equal) best ECE (except for the Song dataset, where it falls short of the LLA ECE by $\mathit{0.1\%}$). It has comparable or better NLL than other methods on all datasets, and often gives an improvement on RMSE. Finally, it is the quickest method, often by a very significant margin, and it does not fail on any datasets, like the other methods do. Note that for the two largest datasets, Protein and Song, we required approximations on the covariance structure of LLA (see (Daxberger et al., 2021)). A detailed explanation of the hyper-parameter tuning method for NUQLS is given in Section C.1.

### v2

@@ 3.2 Theoretical Analysis
$\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})$ $={\bf f}(\widehat{\bm{\theta}},{\bf x})+{{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}\left({\bf w}-{\mathbf{J}}_{\mathcal{X}}\widehat{\bm{\theta}}\right)+\left({{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}-{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x})\right){\bf z}_{0},$ (6)
where ${{\mathbf{K}}}_{{\bf x},\mathcal{X}}\triangleq{{\mathbf{K}}}_{\widehat{\bm{\theta}}}({\bf x},\mathcal{X})\in^{c\times nc}$. Taking ${\bf z}_{0}$ to be a random variable, we form an ensemble of predictors $\{\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\}_{{\bf z}}$, where each $\bm{\theta}^{\star}$ is formed from the projection of a random $\bf z$} onto $\text{Null}({\mathbf{J}}_{\mathcal{X}})$. We require $\bf z$}’s distribution to be symmetric, isotropic, and centered at $\widehat{\bm{\theta}}$, as we do not know a priori which directions contain more information. We take the maximum entropy distribution for a given mean and variance, which is Gaussian [^1] Heavier-tailed distributions matching the above criteria, e.g. logistic distributions, may improve results.. Hence, we let ${\bf z}_{0}\sim\mathcal{N}({\bf 0},\gamma^{2}\mathbf{I})$, for some hyper-parameter $\gamma\in$. The expectation and variance of ˜6, and the distribution of the predictor ${\bf f}(\bm{\theta}^{\star},{\bf x})$, are then
${\mathbb{E}}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\big)=\mathbf{\bm{\mu}}(\widehat{\bm{\theta}},{\bf x})$ $={{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}(\bm{\theta}^{\ddagger}-\widehat{\bm{\theta}})+{\bf f}(\widehat{\bm{\theta}},{\bf x}),$ (7)
$\text{Var}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\big)=\mathbf{\bm{\sigma}}^{2}(\widehat{\bm{\theta}},{\bf x})$ $=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}.$ (8)
${\bf f}(\bm{\theta}^{\star},{\bf x})\overset{\text{approx}}{\sim}$ $\mathcal{N}\big(\bm{\mu}(\widehat{\bm{\theta}},{\bf x}),\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})\big).$ (9)
@@
[…]
@@
Beyond quadratic loss, for a general loss function satisfying the assumptions of Lemma˜3.1, a clear GP posterior interpretation like that in ˜10 may not exist. Nevertheless, we can still derive related insights. If $\widehat{\bm{\theta}}$ is an interpolating solution from initial training, which is common for modern NNs, then as long as ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank, solving ˜3 is equivalent to finding $\bm{\theta}\in^{p}$ such that $\widetilde{{\bf f}}(\bm{\theta},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., find $\bm{\theta}\in^{p}$ such that ${\mathbf{J}}_{\mathcal{X}}(\bm{\theta}-\widehat{\bm{\theta}})={\bf 0}$. So we obtain (9) with $\bm{\mu}(\widehat{\bm{\theta}},{\bf x})={\bf f}(\widehat{\bm{\theta}},{\bf x})$ and $\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}$ as the conditional distribution of ˜10, conditioned on the event ${\bf f}(\bm{\theta}^{\star},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., interpolation with $\ell({\bf f}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})=0$ for $i=1,\ldots,n$.
@@
Drawing samples from the posterior ˜9 by explicitly calculating ˜7 and 8 can be intractable in large-scale settings. Moreover, it can be numerically unstable due to the highly ill-conditioned nature of the NTK matrix [^2] Recall that the condition number of ${{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}$ is the square of that of ${\mathbf{J}}_{\mathcal{X}}$.. However, by combining the above derivations with Theorem˜3.2, we arrive at the key property of Algorithm˜1: it enables efficient sampling from the posterior ˜9.
@@
With the assumptions of Theorem˜3.2, the samples generated by Algorithm˜1 represent draws from the predictive distribution in ˜9.
Hence, we approximate ˜7 and 8 by computing the sample mean and covariance of $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x})\}_{s=1}^{S}$ obtained from Algorithm˜1. By the law of large numbers, the quality of these approximations improves as $S\to\infty$.
@@
Loss functions that do not satisfy the assumption of Lemma˜3.1, such as the cross-entropy loss, may fail to yield a unique representation of ˜4, so the above posterior analysis does not apply. However, our experiments demonstrate that Algorithm˜1 can still generate samples that effectively capture the posterior variance (see Section˜4.4) and posterior mean (see Section˜G.1). In cases where ˜3 lacks a solution, Algorithm˜1 can still be executed by terminating the iterations of (stochastic) GD early. Investigating the distribution of the resulting ensemble and its connection to an explicit posterior remains a potential direction for future research.
4 Experiments
We now empirically demonstrate the result of Theorem˜3.2, as well as compare the performance of our method with alternatives on various regression and classification tasks. Implementation details are given in Appendix˜I. The PyTorch implementation of our experiments is available here. We have also released our method as a package. For additional experimental results, please see Appendix˜H.
4.1 Empirical Convergence
[…]
We demonstrate empirically the convergence of the predictive distribution of NUQLS to that of an NTK-GP. We take a series of MLPs with NTK scaling (Jacot et al., 2018), and train these NNs on normalized Gaussian data (a regression task). Under this setting, the condition number of the resultant NTK is reasonable ($\approx 10^{3}-10^{8}$, compared to $\approx 10^{17}$ on some UCI regression datasets). This allows us to explicitly compute the predictive variance for the NTK-GP for some normalized Gaussian test set. We then also compute the predictive variance using NUQLS, and take the $l_{2}$ norm of the difference between the two variance sets, which we term Squared Error of the Variance (SEV). We plot this against number of epochs of training for NUQLS, and the number of NUQLS realizations $s$. We also plot the average training loss of the linearised networks in the NUQLS ensemble. The values are averaged over $10$ random realizations. The results are displayed in Figure˜2. Note that the SEV cannot be zero, due to the condition number of the NTK; this affects the solution of the linear system $K_{XX}\mathbf{y}=K_{Xx}$ in the computation of the variance for the NTK-GP. However, we see clear convergence of our method to the distribution of an NTK-GP as the ensemble members of NUQLS approach their minima, and as the number of ensemble members increases.
4.2 Toy Regression
We compare the performance of our method on a toy regression problem, taken from (Hernández-Lobato & Adams, 2015) and extended in (Park & Blei, 2024). In Figure˜1, we take $20$ uniformly sampled points in the domain $x\in[-4,-2]\cup[2,4]$, and let $y=x^{3}+\epsilon,~\epsilon\sim\mathcal{N}(0,3^{2})$. A small MLP was trained on these data and used for prediction. We apply VI, SWAG, LA, LLA, DE and NUQLS to the network to find a predictive mean and uncertainty. Close to the training data, i.e. in $4,-2]\cup[2,4]$ we expect low uncertainty; outside of this region, the uncertainty should grow with distance from the training points. VI underestimates, while SWAG and LA overestimate the uncertainty. DE grows more uncertain with distance from the training points, however both NUQLS and the LLA contain the underlying target curve within their confidence intervals. Note that deep ensembles output a heteroskedastic variance term, and were trained on a Gaussian likelihood; in comparison, the variances for LLA and NUQLS were computed post-hoc.
4.3 UCI Regression
@@
Table 1: Comparing performance of NUQLS, DE, LLA and SWAG on UCI regression tasks. NUQLS performs as well as or better than all other methods, while showing a speed up over other methods; this speed up increases with the size of the datasets. LLA-K denotes LLA with a KFAC covariance structure. Reported time for NUQLS, LLA and SWAG includes the training time for the original NN, with the run-time of the post-hoc method given in brackets.
@@
Energy NUQLS $\mathbf{0.047\pm}$$\mathbf{0.006}$ $\mathbf{-2.400\pm}$$\mathbf{0.209}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $\mathbf{8.374}~(\mathbf{0.151})$
DE $218\pm0.0$320.032 $\mathbf{-1.651\pm}$$\mathbf{0.783}$ $\mathbf{0.004\pm}$$\mathbf{0.002}$ $102.244$
LLA $\mathbf{0.048\pm}$$\mathbf{0.006}$ $\mathbf{-2.475\pm}$$\mathbf{0.128}$ $\mathbf{0.004\pm}$$\mathbf{0.004}$ 8.491(0.269)8.491~(0.269)
SWAG $\mathbf{0.058\pm}$$\mathbf{0.015}$ $1.950\pm$$0.158$ $080\pm0.0$110.011 45.306(37.084)45.306~(37.084)
Kin8nm NUQLS $\mathbf{0.252\pm}$$\mathbf{0.005}$ $796\pm0.0$250.025 $\mathbf{0.000\pm}$$\mathbf{0.000}$ $\mathbf{26.570}~(\mathbf{0.264})$
DE $\mathbf{0.252\pm}$$\mathbf{0.006}$ $\mathbf{-0.914\pm}$$\mathbf{0.028}$ $002\pm0.0$010.001 $73.967$
LLA $\mathbf{0.260\pm}$$\mathbf{0.010}$ $783\pm0.0$540.054 $\mathbf{0.001\pm}$$\mathbf{0.001}$ 38.272(11.966)38.272~(11.966)
SWAG $0.457\pm$$0.149$ $0.006\pm$$0.295$ $054\pm0.0$120.012 176.569(150.263)176.569~(150.263)
Protein NUQLS $\mathbf{0.623\pm}$$\mathbf{0.005}$ $\mathbf{0.209\pm}$$\mathbf{0.047}$ $\mathbf{0.002\pm}$$\mathbf{0.000}$ $\mathbf{81.264}~(\mathbf{1.356})$
DE $741\pm0.0$520.052 $\mathbf{0.203\pm}$$\mathbf{0.203}$ $\mathbf{0.011\pm}$$\mathbf{0.020}$ $1014.827$
LLA-K $640\pm0.0$070.007 $458\pm0.0$710.071 $\mathbf{0.002\pm}$$\mathbf{0.000}$ 89.414(9.506)89.414~(9.506)
SWAG $730\pm0.0$440.044 $\mathbf{0.187\pm}$$\mathbf{0.080}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ 548.88(468.972)548.88~(468.972)
@@
In Tables˜1 and 9, we compare NUQLS with DE, LLA and SWAG on a series of UCI regression problems. Mean squared error (MSE) and expected calibration error (ECE) respectively evaluate the predictive and UQ performance, with Gaussian negative log likelihood (NLL) evaluating both. See (Nemani et al., 2023a, $\S 4.11$) for an explanation of ECE. We see that NUQLS consistently has the (equal) best ECE (except for the Song dataset, where it falls short of the LLA ECE by $\mathit{0.1\%}$). It has comparable or better NLL than other methods on all datasets, and often gives an improvement on RMSE. Finally, it is the quickest method, often by a very significant margin, and it does not fail on any datasets, like the other methods do. Note that for the two largest datasets, Protein and Song, we required approximations on the covariance structure of LLA (see (Daxberger et al., 2021)). A detailed explanation of the hyper-parameter tuning method for NUQLS is given in Section˜F.1.